# Harmonic covers of skeleta

### v1

@@ 5. Specialisation of Cartier divisors
Suppose $V$ is any normal Cohen-Macaulay scheme, $\mathscr{D}$ is a cartier divisor and $C$ is a normal curve in $V$ defined over a field $\kappa$. Then $V$ contains no associated points [Liu02, §8.2.15] and so the restriction of Cartier divisors $D|_{E_{\tau}}$ is well-defined. We can therefore define the intersection [^6] One can develop a general refined intersection product of ($\mathbb{Q}$-)Cartier divisors and cycles [Ful84, §20] (lacking intersection products of two arbitrary cycles unless ${X}$ is smooth), but won’t need this. product $\mathscr{D}\cdot C\coloneqq\deg_{\kappa}\mathscr{D}|_{C}\in\mathbb{Z}$, and if $\mathscr{D}$ is only assumed $\mathbb{Q}$-Cartier instead of Cartier we let $\mathscr{D}\cdot C=\frac{1}{e}(e\mathscr{D})\cdot C$ where $e\in\mathbb{Z}_{\geq 1}$ is chosen such that $e\mathscr{D}$ is Cartier. We have the usual rules $(\mathscr{D}+\mathscr{D}^{\prime})\cdot C=\mathscr{D}\cdot C+\mathscr{D}^{\prime}\cdot C$, and $\mathrm{div}\,(f)\cdot C=0$, so that $\mathscr{D}\cdot C$ only depends on the linear equivalence class of $\mathscr{D}$.
We will mainly be interested in the case where $V={X}$ is a toroidal $S$-scheme (which is Cohen-Macaulay by [Kat94, §4.1]), $\kappa=\tilde{k}$ and $C=E_{\tau}$ for some vertical ridge $\tau$ of $\mathcal{D}({X}^{\dagger})$. Note that $\mathscr{D}$ intersects $E_{\tau}$ properly if and only if $F_{\mathscr{D}}|_{\tau}$ is the zero function. Since $\tau$ is vertical, $E_{\tau}$ is a proper $\tilde{k}$-curve (since ${X}$ is assumed proper) and $\mathscr{D}\cdot E_{\tau}$ only depends on the linear equivalence class of $\mathscr{D}$.
We will later need the following outcome of the projection formula (see [Ful84, §20]): let $g:V^{\prime}\to V$ be a proper flat dominant morphism of integral normal $S$-schemes, let $\mathscr{D}$ be a $\mathbb{Q}$-cartier divisor. Suppose $c^{\prime}\in Z_{1}(V^{\prime})$ is a vertical $1$-cycle on $V^{\prime}$ (i.e. a linear combination of curves contained in $V^{\prime}_{s}$). Then (using that $\tilde{k}$ is algebraically closed) we have
@@
For instance if $\mathscr{D}=\mathrm{div}\,(\varpi)$ we have $\mathrm{cyc}\mathscr{D}=[X_{s}]=\sum_{v\in\mathcal{D}({X}^{\dagger})(0)}m_{v}[E_{v}]$ by 5.7. If additionally ${X}$ is $\mathbb{Q}$-factorial then the divisors $[E_{v}]$ for $v\in\mathcal{D}({X}^{\dagger})(0)$ are $\mathbb{Q}$-Cartier, so for every vertical 1-cycle $c\in Z_{1}({X})$ we obtain by the projection formula 5.9 the useful relation
@@
Let $(\sigma,\tau)$ be a facet-ridge pair of a $\mathbb{Z}$-PL complex. Write $m_{\sigma/\tau}$ for the unique generator of $\ker(M_{\sigma}\to M_{\tau})$ such that $\langle m_{\sigma/\tau},\cdot\rangle$ is positive on $\sigma\setminus\tau$. Since $m_{\sigma/\tau}$ kills $M_{\tau}$, the map $\langle m_{\sigma/\tau},\cdot\rangle$ descends to a map $N_{\sigma}/N_{\tau}\to\mathbb{Z}$. Let $n_{\sigma/\tau}$ denote the unique generator of $N_{\sigma}/N_{\tau}$ such that $\langle m_{\sigma/\tau},n_{\sigma/\tau}\rangle>0$. We will also view $n_{\sigma/\tau}$ as a vector of $N_{\tau}$, well-defined up to an element of $N_{\tau}$, and call it the primitive normal vector of $\sigma$ relative to $\tau$.
@@
Let $\mathscr{D}$ be a $\mathbb{Q}$-cartier divisor on ${X}^{\dagger}$. We call the support function $F_{\mathscr{D}}$ harmonic (resp. convex (or subharmonic), resp. strongly convex) if for all ridges $\tau$ of $\mathcal{D}({X}^{\dagger})$ we have $\mathscr{D}\cdot E_{\tau}=0$ (resp. $\mathscr{D}\cdot E_{\tau}\geq 0$, resp. $\mathscr{D}\cdot E_{\tau}>0$). This is for instance the case if $\mathscr{D}$ is numerically trivial (resp. nef, resp. ample). The terminology is justified in Section 6.
@@
Let $\Sigma$ be a $\mathbb{Z}$-PL complex. We define the group of codimension 1 cycles $Z^{1}(\Sigma)$ as the group of formal sums of ridges $[\tau]$, and effective cycles are those with positive coefficients. Now suppose $\phi:\Sigma^{\prime}\to\Sigma$ is a balanced cover of $\mathbb{Z}$-PL complexes. Then we define pullbacks of cycles via the rule $f^{*}[\tau]=\sum_{\tau^{\prime}}\deg_{\tau^{\prime}}\phi[\tau^{\prime}]$ and extending linearly. Note that for any cycle $W\in Z^{1}(\Sigma)$ we have $\deg f^{*}W=\deg f\deg W$.
@@
summing over bounded ridges $\tau$ of $\mathcal{D}({X}^{\dagger})$.
@@
Suppose that $x\in F({X}^{\dagger})$. Recall from 3.9 that $\sigma_{x}$ is empty if and only if $x$ lies in the special fiber ${X}_{s}$. Suppose $x\not\in{X}_{s}$, we also say $x$ is horizontal and write $F({X}^{\dagger})^{\mathrm{hor}}$ for the horizontal points. Note that the generic point of ${X}$ is also considered to be horizontal. By 2.9 it follows that for all $x\in F({X}^{\dagger})^{\mathrm{hor}}$ we have a strict closed immersion $E^{\dagger}_{x}=E_{x}(\mathrm{log}D_{x})\to X^{\dagger}$ where $D_{x}=E_{x}\setminus E_{x}^{\circ}$ and $E^{\dagger}_{x}/S^{\dagger}$ is again a proper toroidal $S^{\dagger}$-scheme 777if $x$ is vertical this is not the case, because if we equip $E_{x}$ with the pullback log-structure then we obtain a so-called hollow log-scheme over the log-point $\tilde{k}$.. In particular we have an associated dual complex $\mathcal{D}(E^{\dagger}_{x})$.
@@
The specialisation map extends to $\overline{\mathcal{D}}(\mathfrak{X}^{\dagger})$ by amalgamating the various specialisation maps for $x\in F(X^{\dagger})^{\mathrm{hor}}$:
(5.28) $\overline{\mathrm{sp}}_{*}\mathscr{D}\coloneqq\bigsqcup_{x\in F({X}^{\dagger})^{\mathrm{hor}}}\overline{\mathrm{sp}}_{*}(\mathscr{D}|_{E_{x}})\in Z^{1}(\overline{\mathcal{D}}(X^{\dagger}))\coloneqq\bigsqcup_{x\in F({X}^{\dagger})^{\mathrm{hor}}}Z^{1}(\mathcal{D}(E_{x}^{\dagger}).$
@@
also see Remark 3.10. Note that since we have removed the divisor $D_{{X}}$ it follows that $m(z)$ is never zero and so $\rho(z)$ is well-defined.
Recall that by Section 3.20 the dual complex $\mathcal{D}({X}^{\dagger})$, and thus $\mathrm{sk}({X}^{\dagger})$ carries a $\mathbb{Z}$-PL structure, and thus by Section 4 also a canonical conformal measure $\mu_{\mathrm{sk}({X}^{\dagger})}$. Volumes will always be taken with respect to this measure. If $x\in F({X}^{\dagger})$ is a log-stratum of ${X}^{\dagger}$, by a slight abuse we also denote by $\sigma_{x}$ the associated polyhedron of $\mathrm{sk}({X}^{\dagger})$. Recall that the lattice of $\mathbb{Z}$-PL functions $M_{\sigma_{x}}$ coincides with the group of Cartier divisors on $\mathscr{O}_{{X},x}$ supported on $D_{{X},x}$.
@@
Note that the previous expression only depends on the image of $f$ in $\mathscr{C}_{{X},\xi(z)}$, and so it follows that the support function $F_{\mathscr{D}}$ factors through the retraction map $\overline{\rho}$.
6. Harmonicity
@@
In this section let ${X}^{\dagger}\to X^{\prime\dagger}$ denote a finite cover of proper $\mathbb{Q}$-factorial toroidal $S$-schemes.
@@
As before let $\Sigma({X}^{\dagger})$ for the cone complex associated to ${X}^{\dagger}$, note that $\Sigma({X}^{\dagger})$ is simplicial by the hypothesis that ${X}$ is $\mathbb{Q}$-factorial, see Remark 5.8. Let $\tau$ be a ridge of $\mathcal{D}({X}^{\dagger})$ and let $C(\tau)$ be the cone generated by $\tau$ in $\Sigma({X}^{\dagger})$. For any ray $r\in C(\tau)(1)$ write $\alpha_{r,t}=-[E_{r}]\cdot E_{\tau}\in\mathbb{Q}$. For any vertex $v\in\tau(0)$ write $\alpha_{v,\tau}=\alpha_{r_{v},\tau}$ where $r_{v}$ is the ray in $\Sigma({X}^{\dagger})$ generated by $v$. We define the bounded multiplicity of $\tau$ as
@@
(6.17) $\Delta(F)\coloneqq\sum_{(\sigma,\tau);\ \tau\text{ bounded }}\partial_{\sigma/\tau}F[\tau],$
summing over facet-ridge pairs $(\sigma,\tau)$ of $\mathcal{D}(X^{\dagger})$ such that $\tau$ is bounded (equivalently $E_{\tau}$ is vertical). We call $F$ harmonic near $x\in\Sigma$ in case for all ridges $\tau$ containing $x$, we have
(6.18) $\sum_{\sigma}\partial_{\sigma/\tau}F=0$
@@
Let $F\in M_{\Sigma}$ where $\Sigma=\mathcal{D}({X}^{\dagger})$ is a dual complex. Then
@@
For comparison, note that the slope formula of [GRW16, §6.8] deals with algebraically closed ground fields and principal divisors.
@@
Let $\phi:(\Sigma^{\prime},\alpha^{\prime})\to(\Sigma,\alpha)$ be a balanced finite cover of $\mathbb{Z}$-PL tropical complexes. Then
@@
Both Theorems 6.20 and 6.22 extend to the compactified skeleta (see 5.31) in the obvious way: we have $\overline{\mathrm{sp}}_{*}\mathscr{D}_{F}=\overline{\Delta}(F)$ where we let $\overline{\Delta}(F)\coloneqq\bigsqcup_{x\in F({X}^{\dagger})^{\mathrm{hor}}}\Delta(F|_{\mathcal{D}(E_{x}^{\dagger})})\in Z^{1}(\overline{\mathcal{D}}({X}^{\dagger}))$, and we have $\phi^{*}\overline{\Delta}(F)=\overline{\Delta}(F\circ\phi)$. We leave the details to the interested reader.
7. Complements on log-canonical sheaves
@@
In this section we recall the calculus of logarithmic differential forms for use in Section 8. Theorem 7.28 gives an elementary proof of a formula comparing canonical sheaves to log-canonical sheaves.
@@
Suppose $f^{\dagger}:X^{\dagger}=(X,\mathscr{M}_{X})\to Y^{\dagger}=(Y,\mathscr{M}_{Y})$ is a morphism of log-schemes. The $\mathscr{O}_{X}$-sheaf $\Omega_{X^{\dagger}/Y^{\dagger}}$ is defined as the universal log-derivation $(\mathrm{d},\mathrm{dlog})$ satisfying the rules $\mathrm{d}\,m=m\,\mathrm{dlog}(m)$ and $\mathrm{dlog}(n)=0$ for all sections $m\in\mathscr{M}_{X}$, $n\in f^{-1}\mathscr{M}_{y}$. In fact $\Omega_{X^{\dagger}/Y^{\dagger}}$ is the quotient of $\Omega_{X/Y}\oplus\left(\mathscr{O}_{X}\otimes\mathscr{M}_{X}^{\mathrm{gp}}\right)$ by the aforementioned relations [Ogu18, Proof of IV.1.2.4]. Note that if $f$ is locally of finite type and the log-structures are coherent, then $\Omega_{X^{\dagger}/Y^{\dagger}}$ is a finitely generated $\mathscr{O}_{X}$-module. Similar as for usual schemes, formation of log-Kähler differentials is compatible with arbitrary base change in the category of log-schemes as well in the categories of fine log-schemes or fs log-schemes [Ogu18, IV.1.2.15].
@@
This follows from [HK94, §2.9.2,§2.9.3] in the fine case and the proof also works in the fs case as we explain now. If in [HK94, 2.9.2] the charts $P,Q$ are chosen fs, then so are $Z_{1}^{\dagger}$ and $Z_{2}^{\dagger}$ since they admit charts by $P$ and $Q\oplus\mathbb{N}^{r}$ for some $r\in\mathbb{N}$. The proof of [HK94, 2.9.3] relies on [Kat89, 4.10], and if in the proof of [Kat89, 4.10] the chart $h:P\to Q$ is fs then so is the monoid $Q^{\prime}=(h^{\mathrm{gp}})^{-1}(P)\subset Q^{\mathrm{gp}}$: indeed, if $q\in Q^{\mathrm{gp}}$ and $q^{n}\in Q^{\prime}$ for some $n\in\mathbb{N}$ then $h(q)^{n}=h(q^{n})\in P$ implies $q\in Q^{\prime}$.
@@
Note that since $h^{\dagger}$ is log-smooth the morphism $h^{*}\Omega_{Y^{\dagger}/S^{\dagger}}\to\Omega_{W^{\dagger}/S^{\dagger}}$ is split injective, implying the same is true for $f^{*}\Omega_{Y^{\dagger}/S^{\dagger}}\to i^{*}\Omega_{W^{\dagger}/S^{\dagger}}$. The conclusion now follows from a diagram chase in the following commuting diagram of exact complexes.
@@
In the case where $f^{\dagger}:X^{\dagger}\to Y^{\dagger}$ is an integral morphism, we can also interprete the proof of Proposition 7.10 via the log-cotangent complex: By [Ols05, 6.9] Olsson’s logarithmic cotangent complex $\mathbb{L}_{X^{\dagger}/Y^{\dagger}}$ associated to $f^{\dagger}$ is concentrated in degrees $[-1,0]$, and is quasi-isomorphic to
@@
and the result follows by examination of the associated long exact sequence. Perhaps there is also an approach using Gabber’s cotangent complex [Ols05, §8], but we do not pursue the details.
@@
Note that the upper horizontal arrow is injective, and therefore by abstract properties of pushouts the lower horizontal arrow is so too.
@@
which clearly factors through the relation $\mathrm{d}g=g\,\mathrm{dlog}(g)$. Note that by construction the residue map is surjective and that $\mathrm{res}([(\omega,h\mathrm{dlog}(g))])=0$ only if $h=gk$ for some $k\in\mathscr{O}_{X,x}$, and in that case the log-form $[(\omega,h\mathrm{dlog}(g))]=[(\omega+k\mathrm{d}g,0)]$ comes from $\Omega_{X,x}$. This settles the claim made above: (7.14) is an exact sequence.
@@
Let us note that later in 7.27 we will consider log-canonical modules in a more general setting.

### v2

@@ 5. Specialisation of Cartier divisors
Suppose $V$ is any normal Cohen-Macaulay scheme, $\mathscr{D}$ is a Cartier divisor and $C$ is a normal curve in $V$ defined over a field $\kappa$. Then $V$ contains no associated points [Liu02, §8.2.15] and so the restriction of Cartier divisors $D|_{C}$ is well-defined [Liu02, §7.1.29]. We can therefore define the intersection [^7] One can develop a general refined intersection theory of ($\mathbb{Q}$-)Cartier divisors and cycles [Ful84, §20], but not intersection products of two arbitrary cycles unless ${X}$ is smooth. We won’t need this. product $\mathscr{D}\cdot C\coloneqq\deg_{\kappa}\mathscr{D}|_{C}\in\mathbb{Z}$, and if $\mathscr{D}$ is only assumed $\mathbb{Q}$-Cartier instead of Cartier we let $\mathscr{D}\cdot C=\frac{1}{e}(e\mathscr{D})\cdot C$ where $e\in\mathbb{Z}_{\geq 1}$ is chosen such that $e\mathscr{D}$ is Cartier. We have the usual rules $(\mathscr{D}+\mathscr{D}^{\prime})\cdot C=\mathscr{D}\cdot C+\mathscr{D}^{\prime}\cdot C$, and $\mathrm{div}\,(f)\cdot C=0$, so that $\mathscr{D}\cdot C$ only depends on the linear equivalence class of $\mathscr{D}$.
We will mainly be interested in the case where $V={X}$ is a toroidal $S$-scheme (which is Cohen-Macaulay by [Kat94, §4.1]), $\kappa=\tilde{k}$ and $C=E_{\tau}$ for some vertical ridge $\tau$ of $\mathcal{D}({X}^{\dagger})$. Note that $\mathscr{D}$ intersects $E_{\tau}$ properly if and only if $F_{\mathscr{D}}|_{\tau}$ is the zero function. Since $\tau$ is vertical and ${X}$ is assumed proper, $E_{\tau}$ is a proper $\tilde{k}$-curve and $\mathscr{D}\cdot E_{\tau}$ only depends on the linear equivalence class of $\mathscr{D}$.
We will later need the following outcome of the projection formula (see [Ful84, §20]): let $g:V^{\prime}\to V$ be a proper flat dominant morphism of integral normal $S$-schemes, let $\mathscr{D}$ be a $\mathbb{Q}$-cartier divisor. Suppose $c^{\prime}\in Z_{1}(V^{\prime})$ is a vertical $1$-cycle on $V^{\prime}$ (i.e. a linear combination of curves contained in $V^{\prime}_{s}$). Then using that $\tilde{k}$ is algebraically closed we have
@@
For instance if $\mathscr{D}=\mathrm{div}\,(\varpi)$ we have $\mathrm{cyc}\mathscr{D}=[X_{s}]=\sum_{v\in\mathcal{D}({X}^{\dagger})(0)}m_{v}[E_{v}]$ by 5.7. If additionally ${X}$ is $\mathbb{Q}$-factorial then the divisors $[E_{v}]$ for $v\in\mathcal{D}({X}^{\dagger})(0)$ are $\mathbb{Q}$-Cartier, so for every vertical 1-cycle $c\in Z_{1}({X})$ we obtain by the projection formula discussed in 5.9 the useful relation
@@
Let $(\sigma,\tau)$ be a facet-ridge pair of a $\mathbb{Z}$-PL complex $\Sigma$. Write $m_{\sigma/\tau}$ for the unique generator of $\ker(M_{\sigma}\to M_{\tau})$ such that $\langle m_{\sigma/\tau},\cdot\rangle$ is positive on $\sigma\setminus\tau$. Since $m_{\sigma/\tau}$ kills $N_{\tau}$, the map $\langle m_{\sigma/\tau},\cdot\rangle$ descends to a map $N_{\sigma}/N_{\tau}\to\mathbb{Z}$. Let $n_{\sigma/\tau}$ denote the unique generator of $N_{\sigma}/N_{\tau}$ such that $\langle m_{\sigma/\tau},n_{\sigma/\tau}\rangle>0$. We will also view $n_{\sigma/\tau}$ as a vector of $N_{\tau}$, well-defined up to an element of $N_{\tau}$, and call it the primitive normal vector of $\sigma$ relative to $\tau$.
@@
Let $\mathscr{D}$ be a $\mathbb{Q}$-cartier divisor on ${X}^{\dagger}$. We call the support function $F_{\mathscr{D}}$ harmonic, resp. convex (or subharmonic) or resp. strongly convex if for all ridges $\tau$ of $\mathcal{D}({X}^{\dagger})$ we have $\mathscr{D}\cdot E_{\tau}=0$, $\mathscr{D}\cdot E_{\tau}\geq 0$ or $\mathscr{D}\cdot E_{\tau}>0$ respectively. This is for instance the case if $\mathscr{D}$ is numerically trivial, nef or ample respectively. The terminology is justified in Section 6.
@@
Let $\Sigma$ be a $\mathbb{Z}$-PL complex. We define the group of codimension 1 cycles $Z^{1}(\Sigma)$ as the group of formal sums of ridges $[\tau]$, and effective cycles are those with positive coefficients. Now suppose $\phi:\Sigma^{\prime}\to\Sigma$ is a balanced cover of $\mathbb{Z}$-PL complexes. Then we define pullbacks of cycles via the rule $f^{*}[\tau]=\sum_{\tau^{\prime}}\deg_{\tau^{\prime}}\phi[\tau^{\prime}]$ summing over ridges $\tau^{\prime}$ above $\tau$, and extending linearly. Note that for any cycle $W\in Z^{1}(\Sigma)$ we have $\deg f^{*}W=\deg f\deg W$.
@@
summing over vertical ridges $\tau$ of $\mathcal{D}({X}^{\dagger})$.
@@
Suppose that $x\in F({X}^{\dagger})$. Recall from 3.9 that $\sigma_{x}$ is empty if and only if $x$ lies in the special fiber ${X}_{s}$. Suppose $x\not\in{X}_{s}$, we also say $x$ is horizontal and write $F({X}^{\dagger})^{\mathrm{hor}}$ for the horizontal points. The generic point of ${X}$ is also considered to be horizontal. By 2.9 it follows that for all $x\in F({X}^{\dagger})^{\mathrm{hor}}$ we have a strict closed immersion $E^{\dagger}_{x}=E_{x}(\mathrm{log}D_{x})\to X^{\dagger}$ where $D_{x}=E_{x}\setminus E_{x}^{\circ}$ and $E^{\dagger}_{x}/S^{\dagger}$ is again a proper toroidal $S^{\dagger}$-scheme 888if $x$ is vertical this is not the case, because if we equip $E_{x}$ with the pullback log-structure then we obtain a so-called hollow log-scheme over the log-point $\tilde{k}$.. In particular we have an associated dual complex $\mathcal{D}(E^{\dagger}_{x})$.
@@
The specialisation map extends to $\overline{\mathcal{D}}({X}^{\dagger})$ by amalgamating the various specialisation maps for $x\in F(X^{\dagger})^{\mathrm{hor}}$:
(5.28) $\overline{\mathrm{sp}}_{*}\mathscr{D}\coloneqq\bigsqcup_{x\in F({X}^{\dagger})^{\mathrm{hor}}}\overline{\mathrm{sp}}_{*}(\mathscr{D}|_{E_{x}})\in Z^{1}(\overline{\mathcal{D}}(X^{\dagger}))\coloneqq\bigsqcup_{x\in F({X}^{\dagger})^{\mathrm{hor}}}Z^{1}(\mathcal{D}(E_{x}^{\dagger})).$
@@
also see Remark 3.10. Since we have removed the divisor $D_{{X}}$ it follows that $m(z)$ is never zero and so $\rho(z)$ is well-defined.
Recall that by Section 3.20 the dual complex $\mathcal{D}({X}^{\dagger})$, and thus $\mathrm{sk}({X}^{\dagger})$ carries a $\mathbb{Z}$-PL structure, and thus by Section 4.8 also a canonical conformal measure $\mu_{\mathrm{sk}({X}^{\dagger})}$. Volumes will always be taken with respect to this measure. If $x\in F({X}^{\dagger})$ is a log-stratum of ${X}^{\dagger}$, by a slight abuse we also denote by $\sigma_{x}$ the associated polyhedron of $\mathrm{sk}({X}^{\dagger})$. Recall that the lattice of $\mathbb{Z}$-PL functions $M_{\sigma_{x}}$ coincides with the group of Cartier divisors on $\mathscr{O}_{{X},x}$ supported on $D_{{X},x}$.
@@
The previous expression only depends on the image of $f$ in $\mathscr{C}_{{X},\xi(z)}$, and so it follows that the support function $F_{\mathscr{D}}$ factors through the retraction map $\overline{\rho}$.
6. Harmonic morphism of $\mathbb{Z}$-PL tropical complexes
@@
In this section let $f:{X}^{\dagger}\to X^{\prime\dagger}$ denote a finite cover of proper $\mathbb{Q}$-factorial toroidal $S$-schemes.
@@
As before let $\Sigma({X}^{\dagger})$ denote the cone complex associated to ${X}^{\dagger}$, note that $\Sigma({X}^{\dagger})$ is simplicial by the hypothesis that ${X}$ is $\mathbb{Q}$-factorial, see Remark 5.8. Let $\tau$ be a ridge of $\mathcal{D}({X}^{\dagger})$ and let $C(\tau)$ be the cone generated by $\tau$ in $\Sigma({X}^{\dagger})$. For any ray $r\in C(\tau)(1)$ write $\alpha_{r,t}=-[E_{r}]\cdot E_{\tau}\in\mathbb{Q}$. For any vertex $v\in\tau(0)$ write $\alpha_{v,\tau}=\alpha_{r_{v},\tau}$ where $r_{v}$ is the ray in $\Sigma({X}^{\dagger})$ generated by $v$. We define the bounded multiplicity of $\tau$ as
@@
(6.17) $\Delta(F)\coloneqq\sum_{(\sigma,\tau);\ \tau\text{ vertical}}\partial_{\sigma/\tau}F[\tau],$
summing over facet-ridge pairs $(\sigma,\tau)$ of $\mathcal{D}(X^{\dagger})$ such that $\tau$ is vertical. We call $F$ harmonic near $x\in\Sigma$ in case for all ridges $\tau$ containing $x$, we have
(6.18) $\sum_{\sigma}\partial_{\sigma/\tau}F=0,$
summing over facets $\sigma$ containing $\tau$.
@@
Let ${X}^{\dagger}$ be a proper $\mathbb{Q}$-factorial toroidal $S$-scheme. Then for any $F\in M_{\mathcal{D}({X}^{\dagger})}$ we have
@@
For comparison we note that the slope formula of [GRW16, §6.8] deals with algebraically closed ground fields and principal divisors.
@@
Let $\phi:(\Sigma^{\prime},\alpha^{\prime})\to(\Sigma,\alpha)$ be a balanced finite cover of $\mathbb{Z}$-PL tropical complexes (Definition 6.12). Then
@@
Both Theorems 6.20 and 6.22 extend to the compactified skeleta (see 5.31) as follows: we have $\overline{\mathrm{sp}}_{*}\mathscr{D}_{F}=\overline{\Delta}(F)$ where we let $\overline{\Delta}(F)\coloneqq\bigsqcup_{x\in F({X}^{\dagger})^{\mathrm{hor}}}\Delta(F|_{\mathcal{D}(E_{x}^{\dagger})})\in Z^{1}(\overline{\mathcal{D}}({X}^{\dagger}))$, and we have $\phi^{*}\overline{\Delta}(F)=\overline{\Delta}(F\circ\phi)$. We leave the details to the interested reader.
7. Complements on determinants of log-differentials
@@
In this section we discuss some calculus of logarithmic differential forms for later use in Section 8. Theorem 7.28 gives an elementary proof of a formula comparing canonical sheaves to log-canonical sheaves.
@@
Suppose $f^{\dagger}:X^{\dagger}=(X,\mathscr{M}_{X})\to Y^{\dagger}=(Y,\mathscr{M}_{Y})$ is a morphism of log-schemes. The $\mathscr{O}_{X}$-sheaf $\Omega_{X^{\dagger}/Y^{\dagger}}$ is defined as the universal log-derivation $(\mathrm{d},\mathrm{dlog})$ satisfying the rules $\mathrm{d}\,m=m\,\mathrm{dlog}(m)$ and $\mathrm{dlog}(n)=0$ for all sections $m\in\mathscr{M}_{X}$, $n\in f^{-1}\mathscr{M}_{y}$. In fact $\Omega_{X^{\dagger}/Y^{\dagger}}$ is the quotient of $\Omega_{X/Y}\oplus\left(\mathscr{O}_{X}\otimes\mathscr{M}_{X}^{\mathrm{gp}}\right)$ by the aforementioned relations [Ogu18, Proof of IV.1.2.4]. If $f$ is locally of finite type and the log-structures are coherent, then $\Omega_{X^{\dagger}/Y^{\dagger}}$ is a finitely generated $\mathscr{O}_{X}$-module. Similar as for usual schemes, formation of log-Kähler differentials is compatible with arbitrary base change in the category of log-schemes as well in the categories of fine log-schemes or fs log-schemes [Ogu18, IV.1.2.15].
@@
This follows from [HK94, §2.9.2,§2.9.3] in the fine case and the proof also works in the fs case as we explain now. If in [HK94, 2.9.2] the charts $P,Q$ are chosen fs, then so are $Z_{1}^{\dagger}$ and $Z_{2}^{\dagger}$ since they admit charts by $P$ and $Q\oplus\mathbb{N}^{r}$ for some $r\in\mathbb{N}$. The proof of [HK94, 2.9.3] relies on [Kat89, 4.10], and if in the proof of [Kat89, 4.10] the chart $h:P\to Q$ is fs then so is the monoid $Q^{\prime}=(h^{\mathrm{gp}})^{-1}(P)\subset Q^{\mathrm{gp}}$: indeed, if $q\in Q^{\mathrm{gp}}$ and $q^{n}\in Q^{\prime}$ for some $n\in\mathbb{N}$ then $h(q)^{n}=h(q^{n})\in P$ implies $q\in Q^{\prime}$. ∎
@@
Since $h^{\dagger}$ is log-smooth the morphism $h^{*}\Omega_{Y^{\dagger}/S^{\dagger}}\to\Omega_{W^{\dagger}/S^{\dagger}}$ is split injective, implying the same is true for $f^{*}\Omega_{Y^{\dagger}/S^{\dagger}}\to i^{*}\Omega_{W^{\dagger}/S^{\dagger}}$. The conclusion now follows from a diagram chase in the following commuting diagram of exact complexes.
@@
In the special case where $f^{\dagger}:X^{\dagger}\to Y^{\dagger}$ is an integral morphism, we can also interpret the proof of Proposition 7.10 via the log-cotangent complex: By [Ols05, 6.9] Olsson’s logarithmic cotangent complex $\mathbb{L}_{X^{\dagger}/Y^{\dagger}}$ associated to $f^{\dagger}$ is concentrated in degrees $[-1,0]$, and is quasi-isomorphic to
@@
and the result follows by examination of the associated long exact sequence. Perhaps there is also an approach using Gabber’s log-cotangent complex [Ols05, §8], but we do not pursue the details.
@@
The upper horizontal arrow is injective, and therefore by abstract properties of pushouts the lower horizontal arrow is so too.
@@
which factors through the relation $\mathrm{d}g=g\,\mathrm{dlog}(g)$. By construction the residue map is surjective and $\mathrm{res}([(\omega,h\mathrm{dlog}(g))])=0$ only if $h=gk$ for some $k\in\mathscr{O}_{X,x}$, in that case the log-form $[(\omega,h\mathrm{dlog}(g))]=[(\omega+k\mathrm{d}g,0)]$ comes from $\Omega_{X,x}$. This settles the claim that (7.14) is an exact sequence.
@@
Later in 7.27 we will consider log-canonical modules in a more general setting.